# On combinatorial bounds for the total Tjurina numbers of certain curves and surfaces with isolated singularities

### v1

@@ Abstract
As an application, we construct a particular family of surfaces in $\mathbb{P}^{3}$ with only isolated singularities having arbitrarily large total Tjurina numbers. This provides new lower bounds for the total Tjurina numbers of certain hypersurfaces that are independent of detailed homological data.
@@ 1 Introduction
The total Tjurina number of a projective hypersurface measures the complexity of its singularities and plays a central role in the interplay between singularity theory, combinatorics, and the homological properties of Jacobian algebras. For plane curves, a classical result of du Plessis and Wall establishes sharp upper bounds on the total Tjurina number in terms of the degree and the minimal degree of Jacobian relations, with equality characterizing free curves [6]. While these upper bounds are rather well understood, see for instance [2], much less is known about effective lower bounds, especially in geometrically constrained situations such as line or conic-line arrangements with prescribed singularities. The aim of this paper is to establish combinatorial lower bounds for the total Tjurina numbers of plane curve arrangements with only ordinary quasi-homogeneous singularities, with a particular focus on free arrangements. This work continues the line of research we initiated recently in [14], where we studied the freeness of conic-line arrangements with ordinary quasi-homogeneous singularties. Our results are motivated by recent developments in the theory of plane curves: by imposing natural geometric conditions, we can observe that free arrangements are highly constrained. We first derive structural inequalities for complex line arrangements with bounded multiplicities, demonstrating that the combinatorics of intersection points strongly restrict possible singular configurations. In particular, Theorem 3.5 shows that there are only finitely many combinatiorial types of supersolvable line arrangements admitting points of multiplicity $\leq 4$. We then consider free conic-line arrangements with ordinary quasi-homogeneous singularities, proving that their total Tjurina numbers grow at least quadratically in the number of components. This constitutes the main result of the paper, see Theorem 4.1.
@@ 2 Preliminaries
Throughout the paper we work over the field of complex numbers $\mathbb{C}$. All varieties are assumed to be reduced and projective. We will follow the notation captured from [1]. Let $S=\mathbb{C}[x,y,z]$ be the graded polynomial ring in three variables $x,y,z$.
@@
If $p\in{\rm Sing}(C)$ is an ordinary quasi-homogeneous singularity, then
@@ 3 Results concerning complex line arrangements with small multiplicities
Combining this with the Hirzebruch-type inequality (1), we get
@@ 4 A lower bound on the total Tjurina numbers of some conic-line arrangements
Recall that the following the naïve combinatorial count holds:
@@
where $f_{i}=\sum_{r\geq 2}r^{i}n_{r}$. Let us come back to (2), we can rewrite it as
@@
We compute the discriminant of the above equation
@@
Finally, we can focus on the total Tjurina number of $\mathcal{CL}$. We have
@@
The obtained lower bound is pretty accurate as the following example shows. Let us recall that the Klein arrangement $\mathcal{K}$ of $21$ lines is a free line arrangement such that $n_{3}=28$ and $n_{4}=21$, see [11, Chapter 6].
@@ 5 An application towards surfaces with only isolated singularities
If $X=\{F(x,y,z,w)=0\}\subset\mathbb{P}^{3}$ us a reduced surface of degree $d$ with at most isolated singularities, the
@@
The fact that $X$ has only isolated singularities follows from straightforward computations of its partial derivatives. In order to find a lower bound on $\tau(X)$ we notice that this is a local problem i.e., we perform computations around each singular point, and we have a naïve Thom-Sebastiani principle for isolated singularities, namely

### v2

@@ Abstract
As an application of these planar results, we construct a family of surfaces in $\mathbb{P}^{3}$ with only isolated singularities and arbitrarily large total Tjurina numbers.This provides new lower bounds for the total Tjurina numbers of certain hypersurfaces that are independent of detailed homological data.
@@ 1 Introduction
The total Tjurina number of a projective hypersurface measures the complexity of its singularities and plays a central role in the interplay between singularity theory, combinatorics, and the homological properties of Jacobian algebras. For plane curves, a classical result of du Plessis and Wall establishes sharp upper bounds on the total Tjurina number in terms of the degree and the minimal degree of Jacobian relations, with equality characterizing free curves [6]. While these upper bounds are rather well understood, see for instance [2], much less is known about effective lower bounds, especially in geometrically constrained situations such as line or conic-line arrangements with prescribed singularities. The aim of this paper is to establish combinatorial lower bounds for the total Tjurina numbers of plane curve arrangements with only ordinary quasi-homogeneous singularities, with a particular focus on free arrangements. This work continues the line of research we initiated recently in [14], where we studied the freeness of conic-line arrangements with ordinary quasi-homogeneous singularities, and shifts the focus toward effective lower bounds for their total Tjurina numbers. Our results are motivated by recent developments in the theory of plane curves: by imposing natural geometric conditions, we can observe that free arrangements are highly constrained. We first derive structural inequalities for complex line arrangements with bounded multiplicities, demonstrating that the combinatorics of intersection points strongly restrict possible singular configurations. In particular, Theorem 3.5 shows that there are only finitely many combinatiorial types of supersolvable line arrangements admitting points of multiplicity $\leq 4$. We then consider free conic-line arrangements with ordinary quasi-homogeneous singularities, proving that their total Tjurina numbers grow at least quadratically in the number of components. This constitutes the main result of the paper, see Theorem 4.1.
@@ 2 Preliminaries
Throughout the paper we work over the field of complex numbers $\mathbb{C}$. All varieties are assumed to be reduced and projective. We follow the notation introduced in [1]. Let $S=\mathbb{C}[x,y,z]$ be the graded polynomial ring in three variables $x,y,z$.
@@
If $p\in{\rm Sing}(C)$ is an ordinary quasi-homogeneous singularity of multiplicity $r$, then
@@ 3 Results concerning complex line arrangements with small multiplicities
Combining this identity with the Hirzebruch-type inequality(1), we obtain
@@
Let $\mathcal{L}\subset\mathbb{P}^{2}$ be a free arrangement of $d\geq 3$ lines such that $m(\mathcal{L})=4$. Then
@@ 4 A lower bound on the total Tjurina numbers of some conic-line arrangements
Recall that the following naïve combinatorial count holds:
@@
where $f_{i}=\sum_{r\geq 2}r^{i}n_{r}$. Let us come back to (2). We can rewrite it as
@@
We compute the discriminant of the above equation in $d_{1}$:
@@
We now turn to estimating the total Tjurina number of $\mathcal{CL}$. We have
@@
The obtained lower bound is quite accurate, as the following example shows. Let us recall that the Klein arrangement $\mathcal{K}$ of $21$ lines is a free line arrangement such that $n_{3}=28$ and $n_{4}=21$, see [11, Chapter 6].
@@ 5 An application towards surfaces with only isolated singularities
If $X=\{F(x,y,z,w)=0\}\subset\mathbb{P}^{3}$ is a reduced surface of degree $d$ with at most isolated singularities, the
@@
The fact that $X$ has only isolated singularities follows from straightforward computations of its partial derivatives. To obtain a lower bound on $\tau(X)$, we note that the problem is local. Hence, we perform computations around each singular point, and we have a naïve Thom-Sebastiani principle for isolated singularities, namely